Solve x^2 + 6x + 9 = 0

Answer by stanbon(75887)

Solve x^2 + 6x + 9 = 0
---
Factor:
(x+3)^2 = 0
Solve:
x+3 = 0
x = -3 with multiplicity 2
=============================
Cheers,
Stan H.
==========